# Emergence of Quantised Representations Isolated to Anisotropic Functions

### v1

@@ 2 Privileged-Plane Projective Method Implementation
Finally, using the vector magnitudes one can produce an angular threshold: $\left\|\vec{v}_{\parallel}\right\|/\left\|\vec{v}\right\|>\epsilon$, where $\theta_{\text{threshold}}=\arccos\epsilon$. Any representations which meet this angular threshold have associated $\vec{w}$ displayed in the PPP method’s plot. This is repeated for all pairwise vectors chosen and all representations chosen.
@@
In this paper, $\epsilon=0.75$ was found to work well and was therefore used in the production of all results. Similar results were found for larger $\epsilon$, but the plots had fewer representations due to the smaller volume, so structures were less definitive in their appearance. Unless otherwise stated, all measurements were taken using the combination-PPP method, using the standard basis. This was found to be beneficial compared to permutation-PPP which could produce statistical artefacts discussed further in App. A.
@@ 3 Experiments and Interpretation
In this section, comparisons between isotropic and anisotropic activation functions are displayed and discussed. The examples shown are an excerpt from a larger variety of results available in App. B. The exact method for training these networks is discussed in App. D.
@@
Figure 1: Displays ten rows of PPP-method’s results, divided into columns. Each row represents an independent autoencoder network trained on the reconstruction of the CIFAR dataset [7]. Hot-spots indicate where the internal latent layer is particularly dense with representations — collated over all samples from the CIFAR training set. The top five rows depict networks which utilise the anisotropic activation function, standard-tanh, whilst the bottom five rows utilise the isotropic activation function, isotropic-tanh. Every other detail is identical otherwise. Each column represents the results of the PPP method at various stages of training. The leftmost column shows a freshly initialised model before training, and moving rightward, the network is progressively trained for up to 125 epochs, as shown in the rightmost column. Figure titles indicate the exact number of epochs trained for. This specific network consists of a latent layer of 18 neurons, with standard (unnormalised) input-output pairs drawn from CIFAR. The dark centres about the origin are attributed to a vanishing volume due to an angular threshold, rather than the absence of representations.
In all cases from Fig. 1, one can observe that the anisotropic row representations form distinct highly-dense, approximately discrete structures, which through training align (horizontal and vertical) or anti-align (diagonal) with the distinguished directions. In contrast, the bottom-five isotropic rows produce more circularly symmetric, continuous representations with no specific alignment tendencies. Furthermore, this example alone is considered sufficiently indicative of a representational inductive bias deriving from functional form choices, thereby confirming this hypothesis. This supports the hypothesis that anisotropic forms are sufficient to cause discrete representational tendency with task-agnostic clustering about distinguished directions.
@@
Firstly, the construction of this specific autoencoder ensured no activation function preceded the latent layer of study, similarly to the spotlight resonance method. This eliminated the possibility that these alignment phenomena are trivially due to the bounding geometry of an activation function, including limit points. The activation functions are only applied after the latent layer of study. Hence, this indicates how the functional form in later layers can exert a representational bias through optimisation, more generally through the network. Thus, the functional form structure biases the parameter trajectories during learning, which then transforms the representations into these discrete structures. Therefore, this represents a non-trivial representational bias due to functional forms confirming the prior prediction [2] in the general case preceding any trivial bounding or neural refraction [2]. Further results also substantiate this.
@@
Additionally, the results suggest that discrete-representational Superposition [1] may arise uniquely in the anisotropic functional form. Several rows, namely 2 through 5, display discrete directions that are either aligned with or anti-aligned with the standard basis, but are sometimes also arranged at more general angles. This is further demonstrated in Fig. 2, which is the same autoencoder but with a greater number of hidden latent layers.
@@
Both are believed to be indicative of a discrete representational structure consistent with descriptions of Superposed features [^3] These are thought to be representational examples of the phenomenon, which differs from the original measurement, which was often concerning the autoencoding network’s parameters. Thus, this potentially provides a sort of dual observation of the phenomenon. Additionally, row 2 demonstrates exclusively anti-aligned representations to the standard basis, which appears to diverge from the classical alignment-first features. [1]. Additionally, it also shows that activation functions can exhert influence backwards through the model, affecting representations in latent layers that precede any activation function in the forward pass. This reframes several Superposition experiments [1], which identifies the phenomenon in such cases.
@@
This may indicate that discrete Superposition phenomena arise due to anisotropic forms, but whether these forms provide a complete description remains debatable — especially for the phenomenon’s dual through parameters, which could still be applicable under continuous representations. Further analysis of individual planes from the PPP method may be necessary to assess Superposition, rather than relying on this more conglomerate plot over all planes. However, this would require many more samples to populate individual planes sufficiently. Additionally, several columns on isotropic networks for early-stage training also indicate directions of high density, but these progressively distribute throughout training. These are believed to be statistical in nature and discussed further in App. A. Overall, it is suggestive that representational Superposition may be a network-adaptive emergent phenomenon predicated upon the imposed task-agnostic anisotropic biases, though a more detailed study would be required.
@@
Generally, anisotropic functional forms appear to produce an increasingly discrete structure in representations as training progresses, beginning with angularly arranged but more diffuse representations and slowly refining into narrow angular beams, consistent with a linear feature hypothesis. This structure appears contingent on the anisotropic functional form, suggesting it carries a quantising inductive bias. Additionally, the anisotropic network’s linear features align about both positive and negative distinguished directions, appearing to respect the sign-flip and permutation (hyperoctahedral symmetry) of the standard-tanh functional form. This is contrasted with the isotropic form examples, which retain a general continuous representational structure which becomes decreasingly discrete through training. Furthermore, hot spots in isotropic representations are more generally angularly distributed and broader, without forming distinct geometries as in the anisotropic cases (such as the high densities along orthogonal directions). This may suggest that the distribution produced in isotropic networks more closely represents the natural structure of the dataset, as it is likely to have more varied angular clusters. Yet, overall, this demonstrates that the representations appear to be influenced by the symmetry structure of the functional forms.
@@
Together, these observed structures are consistent with the sign-flip and permutation (hyperoctahedral) symmetry of the standard basis, which defines the construction of these anisotropic forms. In contrast, the continuous orthogonal functional form yields a smoother, more continuous distribution. This may indicate that the maximal symmetry family is a useful predictor for some representational inductive biases. Comparisons between additional functions derived from each form can substantiate this connection. Furthermore, it suggests that symmetry taxonomies are potentially important to investigate in terms of interpretability, representational capacity, and resultant performance.
@@
This is because the colour map chosen for the PPP method is calculated row-wise, such that all plots on the same row have identical colour map thresholds, enabling better comparison for structure development through training. However, in later layers, the clustering towards these discrete directions is so high that the relative densities in early plots are comparatively negligible and are therefore coloured black by the colour map. This was observed in the number of samples provided by the PPP method. For these plots, the typical number of samples that met the threshold was approximately 2000 for both early-layer anisotropic networks and all isotropic networks. However, for late-stage training in anisotropic networks, this typically increases by two orders of magnitude to around 200,000 samples, indicating that a great majority of representations align with the observed structure. In contrast, isotropic representations remain more evenly distributed throughout the space throughout training.
@@
A final observation is that spontaneous symmetry breaking appears to set distinct directions early during training, and later training only subtly refines these directions. This indicates the strong dependence on the initialisation in networks. Additionally, in the future, this could be explored in terms of whether isotropic networks have a slightly reduced dependence on initialisation, due to momentum enabling a traversal of identically flat connected basins along the symmetry group’s orbit [2].
All these observations remain pertinent to Fig. 2, which presents measurements from a similar, yet deeper, autoencoder model. This model notably produces more complicated discrete structures in anisotropic functional forms, indicating that more nuanced representational Superposition may be occurring due to the increased model depth. Additionally, the bottom five rows corresponding to isotropic functional forms have an increased amount of clustering compared to the previous, more uniform distribution of Fig. 1. This clustering is unaligned with the standard basis, particularly evident in rows 6, 8, and 9; however, close inspection of rows 7 and 10 also reveals misalignment. Together with slight clusterings in Fig. 1, this suggests that it is not due to an imposed structure from functional forms and is indicative of a desirable task-driven clustering.
@@
In the appendix results, there are instances where isotropic experiments display slight hotspots, which sometimes align or anti-align with the standard basis for isotropic networks. This would be surprisingly unlikely given an isotropy’s uniform nature and is instead demonstrated to be a statistical artefact following further controls on experimentation in App. A. Namely, this is believed to be a statistical result due to averages over independent variables for anti-aligned positions, whereas averages over correlated variables for axis-aligned positions. The volumes about the distinguished vectors become overrepresented in the plot, due to the repeated intersection of privileged planes derived from the standard basis. Therefore, averages on these intersections display higher variability, due to the inapplicability of the regression-to-the-mean principle in dependent variables. Hence, under- or over-densities may be statistically more prevalent along standard basis directions, but this effect is typically subtle, although it can be observable. This appendix presents demonstrations of rotated measurement bases for identical isotropic models, which continue to display the phenomenon; therefore, it is deemed highly unlikely that persistent unintended anisotropies exist in implementations or true clusters. Instead, they are attributed to the statistical noise inherent to such conglomerated projection methods. This statistical noise may be skewing row 10 of Fig. 2 to appear slightly axis-aligned; alternatively, this could also be entirely coincidental to the experiment run.
Finally, Fig. 3, demonstrates a similar phenomenon in a comparison between standard Leaky-ReLU and isotropic-Leaky-ReLU functions — details of these functions and further tests can be found in App. B.4. Such results are significant, not only due to their higher prevalence than tanh in contemporary models, but because they analytically differ from tanh in almost every regard, except for the underlying symmetry constructions. Therefore, the demonstration of quantising effects of these activation functions is highly indicative that it is the underlying algebraic symmetries which correspond to such an outcome. This isolates algebraic symmetry as a significant predictor of such phenomena, which is considered impactful for interpretability and explainable AI approaches. Furthermore, each symmetry may correspond to a particular inductive bias, supporting the grouping of functions into taxonomic categories, which may aid in generalising the predicted emergent phenomena.
@@
The appendices present a multitude of additional results for autoencoders of varying widths, depths, datasets, input-output normalisations, latent layer and activation-function dependencies. All of these indicate a similar general tendency towards task-agnostic structure in representations arising from the choice of functional form, particularly the algebraic symmetry involved.
@@ 4 Conclusion
In particular, this structure appears indicative of the maximal symmetry group to which the function algebraically belongs. This is achieved through comparisons of various activation functions with differing maximal symmetries, indicating that symmetry is the primary predictor, rather than the specifics of the chosen function.
Tanh-like activation functions are chosen in two ways. The first is when their form is defined through maximally satisfying a permutation and sign-flip algebraic equivariance of the standard basis, termed the hyperoctahedral group. This function definition results in representational clusters about standard distinguished directions, which accumulate through training, progressively approximating a discrete direction. Whereas, functional forms which maximally abide by an orthogonal algebraic equivariance produce apparent smooth continua of representations. This indicates that the hyperoctahedral discrete group is inducing an approximately-discrete structure to emerge in representation space, implying the functional form is acting as a quantising bias on the representations. Similar is found for algebraic permutation symmetry standard Leaky-ReLU, which also produces a quantising bias, whereas isotropic Leaky-ReLU tends to produce smooth continua too. This is highly indicative of the algebraic symmetry corresponding to the bias, as opposed to function specifics, since Tanh-like and Leaky-ReLU-like functions differ substantially in analytical qualities.
@@
Isotropic primitives are suggested to remove structural biases; this does not then imply that the representations must be isotropic, they merely do not tend to produce form-induced task-agnostic aligned structure to the distinguished directions. However, in the datasets studied, the isotropic networks did tend to form representations which smoothly filled the representation space; consequently, enabling a stronger conviction for the conclusions reached. Effectively, this was an incidentally ideal scenario for determining the validity of the hypotheses. Additionally, it is crucial to reiterate that it is the tendency of models to produce task-agnostic induced structure, which is the subject of study; instances where other arrangements emerge do occur and may require an ensemble of repetitions to clarify.
@@
Consequently, it strongly supports the prediction that choices in activation function form are not benign, but impart unintended inductive biases corresponding to their symmetry properties, which are sufficient in many cases to strongly influence representations. The phenomenon is expected to be systematic for more general anisotropic primitive forms.
@@
The empirical evidence thus far appears highly indicative of the link between algebraic symmetries and inductive biases on representations. If such a finding holds over further analyses and is determined to be generally valid over a broad scope of primitives and models, then one might conjecture it a "Principle/Law for Representational Geometry”: Algebraic Symmetries of Functions Utilised in Models Tend to Propagate Their Geometry Into Emergent Representational Structures Through Training. Despite the evidence of this paper, this general statement remains a tentative conjecture at the current time, requiring substantial further inquiry, and the phrasing likely requires precise refinement and mathematical details.
These results motivate what we tentatively call a conjecture for representational geometry: algebraic symmetries in model primitives tend to propagate into learned representations. We view this as a hypothesis requiring replication across architectures and datasets before any ‘law-like’ status can be justified.
@@
Additionally, combinations of network primitives are expected to exert influence on the representations of larger models, potentially in an interacting, non-trivial hierarchical manner. This is partially demonstrated in the deeper models discussed in the appendices. Such conflating and obscuring influences motivated the first-principles, minimalistic, and ablation-study approach undertaken in this paper, which isolates the cause to individual primitives in otherwise discretising bias-free autoencoders, as opposed to larger models.
@@
This work also corroborates the findings from the SRM study [4] and provides a more fundamental causal framework for its observed alignment phenomenon. Coupled with the findings of Spotlight Resonance Method [4], this work suggests a symmetry causal mechanism for the emergence of several interpretability phenomena, suggesting that they are predicated upon the contemporary anisotropic functional forms. This representational response to anisotropic functional forms appears to explain the downstream emergence of grandmother neurons, generally discrete linear features, axis-alignment, and the representational superposition of discrete features, amongst other phenomena. Hence, axis-alignment and standard grandmother neurons are shown to be phenomena stimulated by design choices, rather than being spontaneous and fundamental to deep learning as a paradigm. Whilst the other phenomena appear at least partially dependent upon such choices. Overall, this strongly supports the emergence of axis-aligned findings that are frequently observed [8, 9] and suggests that structural asymmetry in distributions can emerge, particularly from certain functional forms. This is considered as opposing evidence for representational uniformity [10], at least in terms of autoencoding networks. Together, these results may offer explanatory mechanisms to generalise to further representational geometry and interpretability phenomena.
@@
The isotropic deep learning paper [2] predicts that this phenomenon extends to semanticity. These results demonstrate the formation of approximately discrete representations, which, when combined with the spotlight resonance method’s attribution of semantics to these representations, are considered to validate the semantic prediction; however, an end-to-end analysis could also be undertaken.
@@
This direct mode for the formation of grandmother neurons is at present limited to deep learning and does not establish an empirical link to biological systems. Nevertheless, it could be used to formulate an interdisciplinary hypothesis.
Moreover, the prediction that isotropic functional forms eliminate task-agnostic discretising structure is validated. In the original paper [2], the presence of such structure is suggested to be often pathological; preliminary results support this statement for autoencoding networks. Isotropic formulations are shown to enable the systematic elimination of these biases from a network, as the autoencoders tested featured entirely isotropic primitives, except for the activation functions.
@@
Additionally, this may suggest that interpretable models and AI safety are especially tractable due to anisotropic inductive biases. Thus, design choices between anisotropy, isotropy, or other definitions may allow tuning of models to balance these behaviours with representation capacity and performance, if the pathological nature is further validated. This provides a framework and implementable tunable mode, which earlier work encourages for greater interpretability and safety [1]. However, it would also suggest that representations generated from isotropic implementations currently act in opposition to the assumptions of many interpretability approaches. Heuristically, this indicates that isotropic networks yield more unconstrained representations, which are perhaps more natural or inherent to deep learning structures, as they are free from the human-imposed task-agnostic discretising structure imposed by anisotropy. Yet, it is also often this discretising structure which is made interpretable when understanding AI mechanisms. Hence, bridging such an interpretability-tool gap may be beneficial, and the PPP method offers a starting point.
@@
Overall, this implicates discrete-permutation defined forms of activation functions, as stimulating a task-agnostic alignment about the standard basis, whilst continuous-orthogonal defined forms do not. This indicates that forms carry an inductive bias for representations, which appears to be indicative of the maximal symmetry group structure of the form. Activation functions are demonstrated as a direct causal mechanism for quantised representations, and several downstream emergent phenomena are established through connection to prior work [4]. Hence, this approach has established a causal origin for interpretable structure and several interpretability phenomena.

### v2

@@ 2 Privileged-Plane Projective Method Implementation
Finally, using the vector magnitudes one can produce an angular threshold: $\left\|\vec{v}_{\parallel}\right\|/\left\|\vec{v}\right\|>\epsilon$, where $\theta_{\text{threshold}}=\arccos\epsilon$. Any representations which meet this angular threshold have an associated $\vec{w}$ displayed in the PPP method’s plot. This is repeated for all pairwise vectors chosen and all representations chosen.
@@
In this paper, $\epsilon=0.75$ was found to work well and was therefore used in the production of all results. Similar results were found for larger $\epsilon$, but the plots had fewer representations due to the smaller volume, so structures were less definitive in their appearance. Unless otherwise stated, all measurements were taken using the combination-PPP method, using the standard basis. This was found to be beneficial compared to permutation-PPP, which could produce statistical artefacts discussed further in App. A.
@@ 3 Experiments and Interpretation
In this section, comparisons between isotropic and anisotropic activation functions are displayed and discussed. The examples shown are an excerpt from a larger variety of results available in App. B. The exact method for training these networks is discussed in App. D. A code repository for the PPP method is available at https://github.com/GeorgeBird1/Spotlight-Resonance-Method.
@@
Figure 1: Displays ten rows of PPP-method’s results, divided into columns. Each row represents an independent autoencoder network trained on the reconstruction of the CIFAR dataset Alex (2009). Each column represents the results of the PPP method at various stages of training. The leftmost column shows a freshly initialised model before training, and moving rightward, the network is progressively trained for up to 125 epochs, as shown in the rightmost column. Hot-spots indicate where the internal latent layer is particularly dense with representations — collated over all samples from the CIFAR training set. The top five rows depict networks which utilise the anisotropic activation function, standard-tanh, whilst the bottom five rows utilise the isotropic activation function, isotropic-tanh. Every other detail is identical otherwise. Figure titles indicate the exact number of epochs trained for. This specific network consists of a latent layer of 18 neurons, with standard (unnormalised) input-output pairs drawn from CIFAR. The dark centres about the origin are attributed to a vanishing volume due to an angular threshold, rather than the absence of representations.
In all cases from Fig. 1, one can observe that the anisotropic row representations form distinct highly-dense, approximately discrete structures, which through training align (horizontal and vertical) or anti-align (diagonal) with the distinguished directions. In contrast, the bottom-five isotropic rows produce more circularly symmetric, continuous representations with no specific alignment tendencies. Furthermore, this example alone is considered sufficiently indicative of a representational inductive bias deriving from functional form choices, thereby validating the hypothesis. This supports anisotropic forms being sufficient to cause discrete representational tendency with task-agnostic clustering about distinguished directions.
@@
Firstly, the construction of this specific autoencoder ensured no activation function preceded the latent layer of study, similarly to the spotlight resonance method. This eliminated the possibility that these alignment phenomena are trivially due to the bounding geometry of an activation function, including limit points. The activation functions are only applied after the latent layer of study. Hence, this indicates how the functional forms in later layers can exert a representational bias generally through optimisation throughout the network. Thus, the functional form structure biases the parameter trajectories during learning, which then transforms the representations into these discrete structures. Therefore, this represents a non-trivial representational bias due to functional forms confirming the prior prediction Bird (2025a) in the general case preceding any trivial bounding or neural refraction Bird (2025a). Further results also substantiate this.
@@
Additionally, the results suggest that discrete-representational Superposition Elhage et al. (2022) may arise uniquely in the anisotropic functional form. Several rows, namely two through five, display discrete directions that are either aligned with or anti-aligned with the standard basis, but are sometimes also arranged at more general angles. This is further demonstrated in Fig. 2, which is the same autoencoder but with a greater number of hidden latent layers.
@@
This is believed to be indicative of a discrete representational structure consistent with descriptions of Superposed features [^4] These are thought to be representational examples of the phenomenon, which differs from the original measurement, which was often concerning the autoencoding network’s parameters. Thus, this potentially provides a sort of dual observation of the phenomenon. Additionally, row two demonstrates exclusively anti-aligned representations to the standard basis, which appears to diverge from the classical alignment-first features. Elhage et al. (2022). Additionally, it also shows that activation functions can exert influence backwards through the model, affecting representations in latent layers that precede any activation function in the forward pass. This reframes several Superposition experiments Elhage et al. (2022), which identify the phenomenon in such cases.
@@
This may indicate that discrete Superposition phenomena arise due to anisotropic forms, but whether these forms provide a complete description remains debatable [^5] Especially for the phenomenon’s dual through parameters, which could still be applicable under continuous representations.. Further analysis of individual planes from the PPP method may be necessary to assess further nuances of the Superposition, rather than relying on this more conglomerate plot over all planes. However, this would require many more samples to populate individual planes sufficiently. Additionally, several columns on isotropic networks for early-stage training also indicate directions of high density, but these progressively distribute throughout training. These are believed to be statistical in nature, as optimisation has not had the opportunity to transfer asymmetry in forms to representations yet. This is discussed further in App. A. Overall, it is suggestive that representational Superposition may be a network-adaptive emergent phenomenon predicated upon the imposed task-agnostic anisotropic biases.
@@
Generally, anisotropic functional forms appear to produce an increasingly discrete structure in representations as training progresses, beginning with angularly arranged but more diffuse representations and slowly refining into narrow angular beams. This is consistent with a linear feature hypothesis Elhage et al. (2022). This structure appears contingent on the anisotropic functional form, suggesting it carries a quantising inductive bias. Additionally, the anisotropic network’s linear features align about both positive and negative distinguished directions, appearing to respect the sign-flip and permutation (hyperoctahedral symmetry) of the standard-tanh functional form. This may be indicative of the analytically degenerate orthants produced by such forms. This is contrasted with the isotropic form examples, which retain a general continuous representational structure which becomes decreasingly discrete through training.
Furthermore, hot spots in isotropic representations are more generally angularly distributed and broader, without forming distinct geometries as in the anisotropic cases (such as the high densities along orthogonal directions). This may suggest that the distribution produced in isotropic networks more closely represents the natural structure of the dataset, as it is likely to have more varied angular clusters. Yet, overall, this demonstrates that the representations appear to be influenced by the symmetry structure of the functional forms.
@@
Overall, these observed structures are consistent with the sign-flip and permutation (hyperoctahedral) symmetry of the standard basis, which defines the construction of these anisotropic forms. In contrast, the continuous orthogonal functional form yields a smoother, more continuous distribution. This may indicate that the maximal symmetry family is a useful predictor for some representational inductive biases. Comparisons between additional functions derived from each form can substantiate this connection. Furthermore, it suggests that symmetry taxonomies are potentially crucial to investigate in terms of interpretability, representational capacity, and resultant performance.
@@
This is due to the colour map chosen for the PPP method, which is calculated row-wise, such that all plots on the same row have identical colour map thresholds, enabling better comparison for structure development through training. However, in later layers, the clustering towards these discrete directions is so high that the relative densities in early plots are comparatively negligible and are therefore coloured black by the colour map. This was observed in the number of samples provided by the PPP method. For these plots, the typical number of samples that met the threshold was approximately 2000 for both early-layer anisotropic networks and all isotropic networks. However, for late-stage training in anisotropic networks, this typically increases by two orders of magnitude to around 200,000 samples, indicating that a great majority of representations align with the observed structure. In contrast, isotropic representations remain more evenly distributed throughout the space throughout training.
@@
A final observation is that spontaneous symmetry breaking appears to set distinct directions early during training, and later training only subtly refines these directions. This indicates the strong dependence on the initialisation in networks and supports considering the probabilistic symmetry taxonomisation and its interactions to produce inductive biases Bird (2025a). Additionally, in the future, this could be explored in terms of whether isotropic networks have a slightly reduced dependence on initialisation, due to momentum enabling a traversal of identically flat connected basins along the symmetry group’s orbit Bird (2025a).
All these observations remain pertinent to Fig. 2, which presents measurements from a similar, yet deeper, autoencoder model.
This model notably produces more complicated discrete structures in anisotropic functional forms, indicating that more nuanced representational Superposition may be occurring due to the increased model depth. Additionally, the bottom five rows corresponding to isotropic functional forms have an increased amount of clustering compared to the previous, more uniform distribution of Fig. 1. This clustering is unaligned with the standard basis, particularly evident in rows 6, 8, and 9, with close inspection of rows 7 and 10 also revealing subtler misalignment. Together with slight clusterings in Fig. 1, this suggests that it is not due to an imposed structure from functional forms and is indicative of a desirable task-driven clustering.
@@
In the appendix results, there are instances where isotropic experiments display slight hotspots, which sometimes align or anti-align with the standard basis for isotropic networks. This would be surprisingly unlikely given an isotropy’s uniform nature and is instead demonstrated to be a statistical artefact following further controls on experimentation in App. A. Namely, this is believed to be a statistical result due to averages over independent variables for anti-aligned positions, whereas averages over correlated variables for axis-aligned positions. The volumes about the distinguished vectors become overrepresented in the plot, due to the repeated intersection of privileged planes derived from the standard basis. Therefore, averages on these intersections display higher variability, due to the inapplicability of the regression-to-the-mean principle in dependent variables. Hence, under- or over-densities may be statistically more prevalent along standard basis directions, but this effect is typically subtle, although it can be observable. This appendix presents demonstrations of rotated measurement bases for identical isotropic models, which continue to display the phenomenon; therefore, it is deemed highly unlikely that persistent unintended anisotropies exist in implementations or true clusters. Instead, they are attributed to the statistical noise inherent to such conglomerated projection methods. This statistical noise may be skewing row 10 of Fig. 2 to appear slightly axis-aligned; alternatively, this could also be entirely coincidental to the experiment run. Nevertheless, the strong trend holds that function-derived symmetries act as inductive biases.
Finally, Fig. 3, demonstrates a similar phenomenon in a comparison between standard Leaky-ReLU and isotropic-Leaky-ReLU functions — details of these functions and further tests can be found in App. B.4. Such results are significant, not only due to the ReLU family’s higher prevalence than tanh in contemporary models, but also because they analytically differ from tanh in almost every regard, except for the underlying symmetry constructions. Therefore, the demonstration of quantising effects of these activation functions is highly indicative that it is the underlying algebraic symmetries which correspond to such an outcome. This isolates algebraic symmetry as a significant predictor of such phenomena, which is considered impactful for interpretability and explainable AI approaches. Furthermore, each symmetry may correspond to a particular inductive bias, supporting the grouping of functions into taxonomic categories, which may aid in generalising the predicted emergent phenomena.
@@
The appendices present a multitude of additional results for autoencoders of varying widths, depths, datasets, input-output normalisations, latent layer and activation-function dependencies. All of these indicate a similar general tendency towards task-agnostic structure in representations arising from the choice of functional form, particularly the algebraic symmetry involved. Moreover, they also suggest that geometries in representational superposition may be more likely a result of symmetry than the orthogonal to antipodal to Thomson geometries, which are suggested to balance representation capacity and interference Elhage et al. (2022). This is because hyperoctahedral tanh forms continue to demonstrate orthogonal arrangements first, yet Leaky-ReLU permutation-only forms move immediately to antipodal arrangements for the same task. This would encourage a reconsideration of function-derived symmetries as the primary causal mode for the particular geometries produced, when balancing interference and representational capacity.
@@ 4 Conclusion
In particular, this structure appears indicative of the maximal symmetry group to which the function algebraically belongs. This is achieved through comparisons of various activation functions with differing maximal symmetries, indicating that symmetry is the primary predictor, rather than the specifics of the chosen function. This empirically motivates a thorough investigation of all primitives in terms of their resultant inductive biases when interacting with general networks, previously encouraged in Bird (2025a).
Tanh-like activation functions are chosen in two ways. The first is when their form is defined through maximally satisfying a permutation and sign-flip algebraic equivariance of the standard basis, termed the hyperoctahedral group. This function definition results in representational clusters about standard distinguished directions, which accumulate through training, progressively approximating a discrete direction. Whereas, functional forms which maximally abide by an orthogonal algebraic equivariance produce apparent smooth continua of representations. This indicates that the hyperoctahedral discrete group is inducing an approximately-discrete structure to emerge in representation space, implying the functional form is acting as a quantising bias on the representations. A similar result is found for algebraic permutation symmetry standard Leaky-ReLU, which also produces a quantising bias, whereas isotropic Leaky-ReLU tends to produce smooth continua too. This is highly indicative of the algebraic symmetry corresponding to the bias, as opposed to solely function specifics, since Tanh-like and Leaky-ReLU-like functions differ substantially in analytical qualities. This empirically validates an earlier hypothesis Bird (2025a).
@@
Isotropic primitives are suggested to remove structural biases; this does not then imply that the representations must be isotropic, they merely do not tend to produce form-induced task-agnostic aligned structure to the distinguished directions. However, in the datasets studied, the isotropic networks did tend to form representations which smoothly filled the representation space; consequently, enabling a stronger conviction for the conclusions reached. Effectively, this was an incidentally ideal scenario for determining the validity of the hypotheses. Additionally, it is crucial to consider only the tendency of models to produce task-agnostic induced structure — this is the subject of study. Instances where other arrangements emerge do naturally arise, and this motivated the ensemble plots of repetitions to clarify.
@@
Consequently, the empirical results strongly support the prediction that choices in activation function form are not benign, but impart unintended inductive biases corresponding to their symmetry properties, which are sufficient in many cases to strongly influence representations previously indicated Bird (2025a). The phenomenon is expected to be systematic for more general anisotropic primitive forms.
@@
The empirical evidence thus far appears highly indicative of the link between algebraic symmetries and inductive biases on representations. If such a finding holds over further analyses and is determined to be generally valid over a broad scope of primitives and models, then one might conjecture that it may constitute a principle for representational geometry: Algebraic Symmetries of Functions Utilised in Models Tend to Propagate Their Geometry Into Emergent Representational Structures Through Training. Despite the evidence of this paper, this general statement remains a tentative conjecture at the current time, requiring substantial further inquiry, with the specific phrasing likely requiring precise refinement and mathematical details. For example, this may not be expected to hold where the symmetry can not act on the network, such as in the case of algebraic end-to-end symmetries, which use it to transform representations with the data structure, unless this is reframed as a side-effect of such an inductive bias.
@@
Additionally, combinations of network primitives are expected to exert influence on the representations of larger models, potentially in an interacting, non-trivial hierarchical manner. Understanding this is a pivotal goal of Taxonomic Deep Learning’s approach Bird (2025a). This is partially demonstrated in the deeper models discussed in the appendices. Yet, overall, these potentially conflating and obscuring influences motivated the first-principles, minimalistic, and ablation-study approach undertaken in this paper, which isolates the cause to individual primitives in otherwise discretising bias-free autoencoders, as opposed to larger models.
@@
This work also corroborates the findings from the SRM study Bird (2025b) and provides a more fundamental causal framework for its observed alignment phenomenon. Coupled with the findings of Spotlight Resonance Method Bird (2025b), this work suggests a function-driven, symmetry-based causal mechanism for the emergence of several interpretability phenomena, suggesting that they are predicated upon the contemporary anisotropic functional forms. This representational response to anisotropic functional forms appears to explain the downstream emergence of grandmother neurons, generally discrete linear features, axis-alignment, and the representational superposition of discrete features, amongst other phenomena. Hence, axis-alignment and standard grandmother neurons are shown to be phenomena stimulated by design choices, rather than being spontaneous and fundamental to deep learning as a paradigm. Whilst the other phenomena appear to be strongly dependent upon such choices. Overall, this strongly supports the emergence of axis-aligned findings that are frequently observed Bau et al. (2017); Olah et al. (2019) and suggests that structural asymmetry in distributions can emerge, particularly from certain functional forms. This is considered as opposing evidence for representational uniformity Szegedy et al. (2014)s. Together, these results may offer explanatory mechanisms to generalise to further representational geometry and interpretability phenomena.
@@
The isotropic deep learning paper Bird (2025a) also predicts that this phenomenon extends to semantics. These results demonstrate the formation of approximately discrete representations, which, when combined with the spotlight resonance method’s attribution of semantics to these representations, are considered to validate the semantic prediction too. An end-to-end analysis could also be undertaken in future work.
This direct mode for the formation of grandmother neurons is, at present, limited to deep learning and does not establish an empirical link to biological systems. Nevertheless, it could be used to formulate an interdisciplinary hypothesis.
@@
Moreover, the prediction that isotropic functional forms eliminate task-agnostic discretising structure is validated Bird (2025a). The presence of such structure is also suggested to be often pathological Bird (2025a); preliminary results support this statement for autoencoding networks. Isotropic formulations are shown to enable the systematic elimination of these biases from a network, as the autoencoders tested featured entirely isotropic primitives, except for the activation functions. This establishes isotropies’ inductive bias over a wider variety of primitives; follow-up can extend this for anisotropic networks by selectively reintroducing permutation-derived functions in other primitives.
@@
Additionally, this may suggest that interpretable models and AI safety are especially tractable due to anisotropic inductive biases. Thus, design choices between anisotropy, isotropy, or other definitions may allow tuning of models to balance these behaviours with representation capacity and performance — especially if the pathological nature is further validated. This provides a framework and implementable tunable mode, which earlier work encourages for greater interpretability and safety Elhage et al. (2022). However, it would also suggest that representations generated from isotropic implementations currently act in opposition to the assumptions of many interpretability approaches. Heuristically, this indicates that isotropic networks yield more unconstrained representations, which are perhaps more natural or inherent to deep learning structures, as they are free from the human-imposed task-agnostic discretising structure imposed by anisotropy. Yet, it is also often this discretising structure which is made interpretable when understanding AI mechanisms. Hence, bridging such an interpretability tool gap may be beneficial, and the PPP method offers a starting point.
@@
Overall, this implicates discrete-permutation defined forms of activation functions, as stimulating a task-agnostic alignment about the standard basis, whilst continuous-orthogonal defined forms do not. This indicates that forms carry an inductive bias for representations, which appears to be indicative of the maximal symmetry group structure of the form. Activation functions are demonstrated as a direct causal mechanism for quantised representations, and several downstream emergent phenomena are established through connection to prior work Bird (2025b). Hence, this approach has established a causal origin for interpretable structure and several interpretability phenomena.